# Dialects for CoAP-like Messaging Protocols

### v1

@@ Abstract
Messaging protocols for resource limited systems such as distributed IoT systems are often vulnerable to attacks due to security choices made to conserve resources such as time, memory, or bandwidth. For example, use of secure layers such as DTLS are resource expensive and can sometimes cause service disruption. Protocol dialects are intended as a light weight, modular mechanism to provide selected security guarantees, such as authentication. In this report we study the CoAP messaging protocol and define two attack models formalizing different vulnerabilities. We propose a generic dialect for CoAP messaging. The CoAP protocol, dialect, and attack models are formalized in the rewriting logic system Maude. A number of case studies are reported illustrating vulnerabilities and effects of applying the dialect. We also prove (stuttering) bisimulations between CoAP messaging applications and dialected versions, thus ensuring that dialecting preserves LTL properties (without Next) of CoAP applications.
@@ 1 Introduction
Existing dialect works, for example [sjoholmsierchio-etal-2021netsoft, gogineni-etal-2022arXiv, ren-etal-2023seccom], usually consider protocols running over TCP or other reliable transport. In this case mutations can reasonably be synchronized based on time. In the case of unreliable transport where messages can be delayed, reordered, lost, or repeated, time-based synchronization is problematic. Unreliable transport is mentioned in [GEMM-2023esorics] as one classification parameter, but synchronization mechanisms for this case are not treated.
In this work we study dialects for protocols running over unreliable transport, using the CoAP messaging protocol [RFC7252] as prototypical example. We investigate the use of counters, analogous to those used to prevent replay in security protocols. for synchronizing. We consider two methods of evaluating the CoAP specification and dialect wrapper. One is based on the CoAP vulnerabilities discussed in [coap-attacks] and the other is based on properties a CoAP enabled endpoint could assume when running over an unreliable transport layer such as UDP.
For the vulnerability case, we formalize the attack scenarios presented in [coap-attacks] as motivation for strengthening the specification. These scenarios consider an attacker that can only block/drop or delay messages. We use reachability analysis to look for attacks. In some cases the Maude endpoints exhibit the expected vulnerability. in some cases because the Maude specification implements some of the stronger mechanisms, the attack fails, but there are related attacks. In some cases an attacker that can redirect messages can realize the attack, but dialecting mitigates such attacks.
@@
  • Reachability analysis demonstrating CoAP vulnerabilities identified in [coap-attacks] (Appendix 0.B). These require an active attacker.
@@
Section 2 provides background information on CoAP, dialects, and rewriting logic/Maude. Section 3 describes the Maude specification of CoAP messaging. Resource limited attack models are presented in Section 4 with examples illustrating a variety of attacks presented in Section 5. In Section 6 we present the abstract mutation functions of the proposed dialect scheme. The Maude specification of the CoAP dialect is given in Section 7. In Section 8 we discuss the properties of dialects such as the CoAP dialect. The relation to [GEMM-2023esorics] is discussed in Section 9. Concluding remarks are given in Section 10.
In Appendix 0.A we spell out key properties of CoAP system executions, including characterizing the minimal time elapse function. In Appendix 0.B we specify the scenarios illustrating CoAP vulnerabilities listed in [coap-attacks] and do reachability analysis to determine which vulnerabilities are realized by the Maude specification. We also analyze the dialected versions of each scenario. In Appendix 0.C we prove two bisimulation theorems for dialected CoAP systems, verifying relations proposed in Section 8.
@@ 2.1 The CoAP protocol
The Constrained Application Protocol (CoAP) [RFC7252, coap-attacks] is an HTTP-like client-server Protocol for use by resource-constrained devices (e.g. low power) and networks (low bandwidth, lossy). CoAP provides a request/response RESTFUL interaction model between application endpoints. Servers hold resources that can be generated, updated, accessed by clients. An endpoint can be a client, a server, or both. A design goal is to keep message overhead small, to limit need for fragmentation. Note that server resources may be connected to actuators having effects on the environment.
@@ 2.2 Dialects
The following are some examples of dialects and their application. Dialects to improve SDN (Software defined network) security are defined and evaluated in [sjoholmsierchio-2019nps, sjoholmsierchio-etal-2021netsoft]. Two transformations based on HMAC over the message data are proposed. One places the HMAC in the session id slot. This limits the HMAC size and thus the hash key is changed every second. For messages that use this slot, a 512-bit HMAC tag is appended to data packets. Sequence numbers are included in the hashed content to prevent replay. Keys are generated from pre-shared symmetric keys, delivered out-of-band. Dialects are implemented using proxies inserted into the communication path to carry out the message transformations. One goal is to protect against TLS downgrade attacks by providing authentication for initial TLS communications.
Dialects for FTP and MQTT protocols using bit shuffling or packet splitting are studied in [gogineni-etal-2022arXiv]. A self-synchronization mechanisms is propsed. In [gogineni-etal-2021site] the authors propose a Deep Neural Network mechanism for synchronization. A dialect is implemented by (conceptually) inserting a separate process between the protocol and the network.
Dialecting is proposed in [ren-etal-2023seccom] to mitigate known exploitable vulnerabilities in IoT devices deployed in large numbers thus enabling massive attacks. The idea is to transform functions that construct and parse message data structures to construct and parse mutated data structures. The message constructor/parser functions are identified automatically and message mutations are typically per message field. Specific mutations are designed using testing. Dialects are implemented by modifying the identified message constructor and parser functions of the protocol implementation. The method is evaluated on implementations of MQTT, LwM2M, and DDS. In the experiments, the mutations successfully blocked reproducible attacks corresponding to a selected set of CVEs reported for the protocols
The above works propose dialects and evaluate them experimentally. In [GEMM-2023esorics] a formal framework for specifying and applying dialects is proposed. The ideas are illustrated using the MQTT protocol. This work and its relation to the current work is discussed in detail in section 9.
@@ 2.3 Rewriting logic and Maude
Maude [clavel-etal-07maudebook] is a language and tool implementing rewriting logic. Maude can be downloaded from the Maude website [maudeweb]. Rewrite theories are represented by Maude modules. Functional modules (declared using the keyword fmod at the beginning and endfm at the end) represent equational theories (with no rules). System modules (declared using the keyword mod at the beginning and endm at the end) represent rewrite theories with rules. We introduce the main Maude syntax elements with examples from the CoAP specification (discussed in detail in Section 3). Sorts and their ordering are declared using the keywords sort and subsort. For example, the following declares two sorts Msg (the message sort) and MsgS (multisets of messages) with Msg (singleton sets) being a subsort of MsgS.
@@ 3.1 Data types
The specification of the message data type is a direct formalization of the (abstract) structure of messages as specified in [RFC7252]. Messages (sort Msg) are constructed by the operation
@@ 3.2 Rules
The rules for handling requests are specified in [RFC7252] section 5.8. The function rcvRequest implements these rules. It first checks whether it has already sent a response to this request using matchingRsp(devatts,msg). If so, if the request is confirmable, the ACK is resent, otherwise the message is ignored. If the request is new the method is computed from the code and the appropriate method specific function is called to process the message. As an example, if the method is "GET", the function rcvGet is called. This function sends a single message in response:
@@
For managing the passing of time, we follow the Real-time Maude approach [olveczky08tacas]. The idea is that there is an earliest time at which some instantaneous rule will be enabled to fire, called the minimal time elapse (mte). If the mte is zero then the enabled rules should be applied until there are no more. If the mte is greater than zero then time passes by this amount, reaching a situation where some rules can fire at the current time (or the state is terminal). It is the job of the specification designer to ensure that only a finitely many rules can fire before time must pass. In the CoAP model, mte is greater than zero if all delays of messages in the network, or an w4Ack attribute, are greater than zero, or a message send is pending but the sndCtr is non-zero. The rule labelled tick formalizes the above. Here nz has sort NzNat (non-zero natural number). The function passTime decrements each message delay in the network and w4Ack attributes and each sndCtr by nz. It also decrements delays of messages in rspSntD attributes, modeling the decrease in lifetime of the identifier and token.
@@ 3.4 CoAP Scenario Specification
A scenario is specified by an intial system configuration. This together with the CoAP rewrite theory determines a set of execution traces and the reachable states of the scenario. To support systematic testing of CoAP messaging and analysis of attacks we defined two functions to generate initial configurations for scenarios tCS2C (test Client Server 2 endpoints) and tCS3C (test Client Server 3 endpoints). These are used in the experiments validating the vulnerabilities discussed in [coap-attacks], described in Appendix 0.B. They were also used in experiments testing various CoAP messaging patterns and effects of dialecting.
@@ 4 Attack Models
We consider two main classes of attacker: active and reactive. The active attacker models the attacker envisioned in [coap-attacks]. Using the formal representation of an active attacker we replicated versions of the attacks postulated in [coap-attacks] (see Appendix 0.B). These attack capabilities (except for the rerouting extension, A1, see below) simply amplify the effects of an unreliable transport, dialecting can not mitigate such attacks, and applications need to deal with the unreliability of the network independently of attackers.
@@ 6 Dialect functions
with $\mathtt{content}\neq\mathtt{content1}$ and $\mathtt{content1}$ is recognized as ill-formed. Alternately, $f_{2}$ could directly return an indication of failure, but only if the input is a correctly transformed content.
@@ 7.1 Dialect messages and configurations
As described in [GEMM-2023esorics] we represent a dialect wrapper by a meta-agent that has the same identifier as the wrapped agent, a conf attribute that encapsulates the base agent and its network stub, and additional attributes for managing dialect transformations.
@@ 9 Related Work
The work most closely related to the present work is [GEMM-2023esorics], which we will refer to as ESORICS23 in the following. Complimentary work on dialects is discussed briefly in Section 2. We refer to ESORICS23 for a more complete review of previous work on dialects.
@@
The ESORICS23 attack model starts with two models proposed in [ren-etal-2023seccom]: the off-path attacker that can send and receive messages, but can not see messages exchanged between its targets; and the on-path attacker with all the capabilities of the off-path attacker, plus the ability to observe communications between honest endpoints. ESORICS23 adds a third model: active on-path attacker with the capabilities of the on-path attacker plus the ability to block, redirect, and alter message content. A refinement is also introduced–(m, n, t) attacker that in time t can observe m messages and block n messages (and send/receive, block, …). The goal of a dialect is low probability that an “attacked” message will be accepted by intended receiver.
In this paper we define two attack models for CoAP messaging systems: active and reactive. In either case, an attacker is allowed a bounded set of actions from a set of capabilities (depending on the model). The bound on number of actions is an alternative to the bounds in the (m,n,t) attack model as a means to limit the attacker strength. The active attacker is based on vulnerabilities identified in [coap-attacks]. The capabilities are: block/drop, delay, replay, or redirect messages. Our active attack model differs from the active on-path model in that attackers don’t receive or create messages. Like [coap-attacks] we assume communications may be encrypted, so focuses on attacks that don’t involve parsing messages – all the attacker learns from a message is the target, source and a content blob.
The ESORICS23 paper is focused on understanding the structure of dialects and attack models and gives some examples and informal statements about the security guarantees of lingos and dialects. In the present paper we formally specify an attack model and carry out reachability analyses concerning specific undesired states of simple applications using the CoAP protocol (inspired by [coap-attacks]). We also prove properties relating executions of dialected and undialected scenarios with and without attacks.
@@ Appendix 0.A Characterization of reachable CoAP system configurations
As for real-time Maude specifications [olveczky08tacas] the rewrite rules for the CoAP specification (and it dialected form) are either instantaneous rules or the tick rule that models passing of time. The minimum time to elapse function, mte(sys), determines when time can pass. The following propositions give essential properties of reachable CoAP and dialected CoAP system configurations according to the value mte(sys).
@@ Appendix 0.B CoAP vulnerabilities
The paper [coap-attacks] motivates an extension to the CoAP standard by describing several vulnerabilities where attacks do not depend on being able to see message content, i.e. these attacks can in principle be carried out against CoAP running over DTLS or other transport providing encryption.
@@
This scenario illustrates the ability of an attacker to block a single message. How the attacker recognizes the intended target is not discussed in [coap-attacks]. Th scenario is adapted from Figures 1,2 of [coap-attacks].
@@
This scenario is based on figure 3 of [coap-attacks].
@@
This scenario is based on Figure 4 of [coap-attacks]. Here the door is initially locked. The client requests unlock, possibly carries out some other interactions, and then requests lock. The attacker captures the unlock request. Since the unlock request is confirmable, the client resends it and the server sends a response to the resent request. The following lock request is processed as expected. Then the attacker releases the captured unlock request and drops the response. Thus the client believes the door is locked when in fact it is unlocked.
@@
This scenario is based on Figure 5 of [coap-attacks]. In this attack the client sends an unlock request, the attacker captures the response, since the resquest is non-confirmable the client doesn’t worry about the lack of response. Later the client sends a lock request that is dropped by the attacker, who releases the unlock response to trick the client into thinking the lock request succeeded, when in fact the door remains unlocked.

### v2

@@ Abstract
Messaging protocols for resource limited systems such as distributed IoT systems are often vulnerable to attacks due to security choices made to conserve resources such as time, memory, or bandwidth. For example, use of secure layers such as DTLS are resource expensive and can sometimes cause service disruption. Protocol dialects are intended as a light weight, modular mechanism to provide selected security guarantees, such as authentication. In this report we study the CoAP messaging protocol and define two attack models formalizing different vulnerabilities. We propose a generic dialect for CoAP messaging. The CoAP protocol, dialect, and attack models are formalized in the rewriting logic system Maude. A number of case studies are reported illustrating vulnerabilities and effects of applying the dialect. We also prove (stuttering) bisimulations between CoAP messaging applications and dialected versions, thus ensuring that dialecting preserves LTL properties (without Next) of CoAP applications. To support search for attacks in complex messaging situations we specify a simple application layer to drive the CoAP messaging and generalize the attack model to support a form of symbolic search for attacks. Two case studies are presented to illustrate the more general attack search.
@@ 1 Introduction
Existing dialect works, for example [11, 3, 8], usually consider protocols running over TCP or other reliable transport. In this case mutations can reasonably be synchronized based on time. In the case of unreliable transport where messages can be delayed, reordered, lost, or repeated, time-based synchronization is problematic. Unreliable transport is mentioned in [2] as one classification parameter, but synchronization mechanisms for this case are not treated.
In this work we study dialects for protocols running over unreliable transport, using the CoAP messaging protocol [9] as prototypical example. We investigate the use of counters, analogous to those used to prevent replay in security protocols. for synchronizing. We consider two methods of evaluating the CoAP specification and dialect wrapper. One is based on the CoAP vulnerabilities discussed in [4] and the other is based on properties a CoAP enabled endpoint could assume when running over an unreliable transport layer such as UDP.
For the vulnerability case, we formalize the attack scenarios presented in [4] as motivation for strengthening the specification. These scenarios consider an attacker that can only block/drop or delay messages. We use reachability analysis to look for attacks. In some cases the Maude endpoints exhibit the expected vulnerability. in some cases because the Maude specification implements some of the stronger mechanisms, the attack fails, but there are related attacks. In some cases an attacker that can redirect messages can realize the attack, but dialecting mitigates such attacks.
@@
  • Reachability analysis demonstrating CoAP vulnerabilities identified in [4] (Appendix 0.B). These require an active attacker.
@@
  • A simple application layer is specified to drive complex messaging scenarios and basic properties defined from which to build invariant specifications. The reactive attack model is generalized to support a form of symbolic search for attacks. (Section 0.D).
@@
Section 2 provides background information on CoAP, dialects, and rewriting logic/Maude. Section 3 describes the Maude specification of CoAP messaging. Resource limited attack models are presented in Section 4 with examples illustrating a variety of attacks presented in Section 5. In Section 6 we present the abstract mutation functions of the proposed dialect scheme. The Maude specification of the CoAP dialect is given in Section 7. In Section 8 we discuss the properties of dialects such as the CoAP dialect. The relation to [2] is discussed in Section 9. Concluding remarks are given in Section 10.
In Appendix 0.A we spell out key properties of CoAP system executions, including characterizing the minimal time elapse function. In Appendix 0.B we specify the scenarios illustrating CoAP vulnerabilities listed in [4] and do reachability analysis to determine which vulnerabilities are realized by the Maude specification. We also analyze the dialected versions of each scenario. In Appendix 0.C we prove two bisimulation theorems for dialected CoAP systems, verifying relations proposed in Section 8. The application layer, generalized attack model and associated case studies are presented in Appendix 0.D.
The Maude specification and case studies can be found at https://github.com/SRI-CSL/VCPublic in the folder CoAPDialect.
@@ 2.1 The CoAP protocol
The Constrained Application Protocol (CoAP) [9, 4] is an HTTP-like client-server Protocol for use by resource-constrained devices (e.g. low power) and networks (low bandwidth, lossy). CoAP provides a request/response RESTFUL interaction model between application endpoints. Servers hold resources that can be generated, updated, accessed by clients. An endpoint can be a client, a server, or both. A design goal is to keep message overhead small, to limit need for fragmentation. Note that server resources may be connected to actuators having effects on the environment.
@@ 2.2 Dialects
The following are some examples of dialects and their application. Dialects to improve SDN (Software defined network) security are defined and evaluated in [10, 11]. Two transformations based on HMAC over the message data are proposed. One places the HMAC in the session id slot. This limits the HMAC size and thus the hash key is changed every second. For messages that use this slot, a 512-bit HMAC tag is appended to data packets. Sequence numbers are included in the hashed content to prevent replay. Keys are generated from pre-shared symmetric keys, delivered out-of-band. Dialects are implemented using proxies inserted into the communication path to carry out the message transformations. One goal is to protect against TLS downgrade attacks by providing authentication for initial TLS communications.
Dialects for FTP and MQTT protocols using bit shuffling or packet splitting are studied in [3]. A self-synchronization mechanisms is propsed. In [5] the authors propose a Deep Neural Network mechanism for synchronization. A dialect is implemented by (conceptually) inserting a separate process between the protocol and the network.
Dialecting is proposed in [8] to mitigate known exploitable vulnerabilities in IoT devices deployed in large numbers thus enabling massive attacks. The idea is to transform functions that construct and parse message data structures to construct and parse mutated data structures. The message constructor/parser functions are identified automatically and message mutations are typically per message field. Specific mutations are designed using testing. Dialects are implemented by modifying the identified message constructor and parser functions of the protocol implementation. The method is evaluated on implementations of MQTT, LwM2M, and DDS. In the experiments, the mutations successfully blocked reproducible attacks corresponding to a selected set of CVEs reported for the protocols
The above works propose dialects and evaluate them experimentally. In [2] a formal framework for specifying and applying dialects is proposed. The ideas are illustrated using the MQTT protocol. This work and its relation to the current work is discussed in detail in section 9.
@@ 2.3 Rewriting logic and Maude
Maude [1] is a language and tool implementing rewriting logic. Maude can be downloaded from the Maude website [12]. Rewrite theories are represented by Maude modules. Functional modules (declared using the keyword fmod at the beginning and endfm at the end) represent equational theories (with no rules). System modules (declared using the keyword mod at the beginning and endm at the end) represent rewrite theories with rules. We introduce the main Maude syntax elements with examples from the CoAP specification (discussed in detail in Section 3). Sorts and their ordering are declared using the keywords sort and subsort. For example, the following declares two sorts Msg (the message sort) and MsgS (multisets of messages) with Msg (singleton sets) being a subsort of MsgS.
@@ 3.1 Data types
The specification of the message data type is a direct formalization of the (abstract) structure of messages as specified in [9]. Messages (sort Msg) are constructed by the operation
@@ 3.2 Rules
The rules for handling requests are specified in [9] section 5.8. The function rcvRequest implements these rules. It first checks whether it has already sent a response to this request using matchingRsp(devatts,msg). If so, if the request is confirmable, the ACK is resent, otherwise the message is ignored. If the request is new the method is computed from the code and the appropriate method specific function is called to process the message. As an example, if the method is "GET", the function rcvGet is called. This function sends a single message in response:
@@
For managing the passing of time, we follow the Real-time Maude approach [7]. The idea is that there is an earliest time at which some instantaneous rule will be enabled to fire, called the minimal time elapse (mte). If the mte is zero then the enabled rules should be applied until there are no more. If the mte is greater than zero then time passes by this amount, reaching a situation where some rules can fire at the current time (or the state is terminal). It is the job of the specification designer to ensure that only a finitely many rules can fire before time must pass. In the CoAP model, mte is greater than zero if all delays of messages in the network, or an w4Ack attribute, are greater than zero, or a message send is pending but the sndCtr is non-zero. The rule labelled tick formalizes the above. Here nz has sort NzNat (non-zero natural number). The function passTime decrements each message delay in the network and w4Ack attributes and each sndCtr by nz. It also decrements delays of messages in rspSntD attributes, modeling the decrease in lifetime of the identifier and token.
@@ 3.4 CoAP Scenario Specification
A scenario is specified by an intial system configuration. This together with the CoAP rewrite theory determines a set of execution traces and the reachable states of the scenario. To support systematic testing of CoAP messaging and analysis of attacks we defined two functions to generate initial configurations for scenarios tCS2C (test Client Server 2 endpoints) and tCS3C (test Client Server 3 endpoints). These are used in the experiments validating the vulnerabilities discussed in [4], described in Appendix 0.B. They were also used in experiments testing various CoAP messaging patterns and effects of dialecting.
@@ 4 Attack Models
We consider two main classes of attacker: active and reactive. The active attacker models the attacker envisioned in [4]. Using the formal representation of an active attacker we replicated versions of the attacks postulated in [4] (see Appendix 0.B). These attack capabilities (except for the rerouting extension, A1, see below) simply amplify the effects of an unreliable transport, dialecting can not mitigate such attacks, and applications need to deal with the unreliability of the network independently of attackers.
@@ 6 Dialect functions
with $\mathtt{content}\neq\mathtt{content1}$ and $\mathtt{content1}$ is recognized as ill-formed. Alternately, $f_{2}$ could directly return an indication of failure, but only if the input is not a correctly transformed content.
@@ 7.1 Dialect messages and configurations
As described in [2] we represent a dialect wrapper by a meta-agent that has the same identifier as the wrapped agent, a conf attribute that encapsulates the base agent and its network stub, and additional attributes for managing dialect transformations.
@@ 9 Related Work
The work most closely related to the present work is [2], which we will refer to as ESORICS23 in the following. Complimentary work on dialects is discussed briefly in Section 2. We refer to ESORICS23 for a more complete review of previous work on dialects.
@@
The ESORICS23 attack model starts with two models proposed in [8]: the off-path attacker that can send and receive messages, but can not see messages exchanged between its targets; and the on-path attacker with all the capabilities of the off-path attacker, plus the ability to observe communications between honest endpoints. ESORICS23 adds a third model: active on-path attacker with the capabilities of the on-path attacker plus the ability to block, redirect, and alter message content. A refinement is also introduced–(m, n, t) attacker that in time t can observe m messages and block n messages (and send/receive, block, …). The goal of a dialect is low probability that an “attacked” message will be accepted by intended receiver.
In this paper we define two attack models for CoAP messaging systems: active and reactive. In either case, an attacker is allowed a bounded set of actions from a set of capabilities (depending on the model). The bound on number of actions is an alternative to the bounds in the (m,n,t) attack model as a means to limit the attacker strength. The active attacker is based on vulnerabilities identified in [4]. The capabilities are: block/drop, delay, replay, or redirect messages. Our active attack model differs from the active on-path model in that attackers don’t receive or create messages. Like [4] we assume communications may be encrypted, so focuses on attacks that don’t involve parsing messages – all the attacker learns from a message is the target, source and a content blob.
The ESORICS23 paper is focused on understanding the structure of dialects and attack models and gives some examples and informal statements about the security guarantees of lingos and dialects. In the present paper we formally specify an attack model and carry out reachability analyses concerning specific undesired states of simple applications using the CoAP protocol (inspired by [4]). We also prove properties relating executions of dialected and undialected scenarios with and without attacks.
@@ Appendix 0.A Characterization of reachable CoAP system configurations
As for real-time Maude specifications [7] the rewrite rules for the CoAP specification (and it dialected form) are either instantaneous rules or the tick rule that models passing of time. The minimum time to elapse function, mte(sys), determines when time can pass. The following propositions give essential properties of reachable CoAP and dialected CoAP system configurations according to the value mte(sys).
@@ Appendix 0.B CoAP vulnerabilities
The paper [4] motivates an extension to the CoAP standard by describing several vulnerabilities where attacks do not depend on being able to see message content, i.e. these attacks can in principle be carried out against CoAP running over DTLS or other transport providing encryption.
@@
This scenario illustrates the ability of an attacker to block a single message. How the attacker recognizes the intended target is not discussed in [4]. Th scenario is adapted from Figures 1,2 of [4].
@@
This scenario is based on figure 3 of [4].
@@
This scenario is based on Figure 4 of [4]. Here the door is initially locked. The client requests unlock, possibly carries out some other interactions, and then requests lock. The attacker captures the unlock request. Since the unlock request is confirmable, the client resends it and the server sends a response to the resent request. The following lock request is processed as expected. Then the attacker releases the captured unlock request and drops the response. Thus the client believes the door is locked when in fact it is unlocked.
@@
This scenario is based on Figure 5 of [4]. In this attack the client sends an unlock request, the attacker captures the response, since the resquest is non-confirmable the client doesn’t worry about the lack of response. Later the client sends a lock request that is dropped by the attacker, who releases the unlock response to trick the client into thinking the lock request succeeded, when in fact the door remains unlocked.